OpenStudy (anonymous):

how to find general solution of: y'''' + ky = 0 someone please help its urgent..

OpenStudy (anonymous):

alternatively set \(\large y(x)=e^{rx}\) then \(r^4-k=0\)

OpenStudy (anonymous):

r u dumb?? do u even know what y'''' means? thi is not how u find roots of an equation i know what the roots are.. i just dont know the way to find it

OpenStudy (anonymous):

if \(k\ge 0\)then \((r^2-k)(r^2+k)=0\) \(r=\pm \sqrt{k}\) and \(r=\pm i\sqrt{k}\) \(y(x)=c_1e^{-\sqrt{k}x}+c_2e^{\sqrt{k}x}+c_4\cos \sqrt{kx}+c_5\sin \sqrt{kx}\)